OpenStudy (anonymous):

How much would $100 invested at 6% interest compounded annually be worth after 20 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

339.96

OpenStudy (anonymous):

it was 320.71

OpenStudy (anonymous):

A = P (1 + R/100)^n here P = $ 100 R/100 = 6/100 = 0.06 n = 20 years so A = 100(1 + 0.06)^20 = 100 (1.06)^20 = 100*3.207 = $320.70